# LoSiA: Efficient High-Rank Fine-Tuning via Subnet Localization and Optimization

### v3

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA, significantly reduce the number of trainable parameters by introducing low-rank decomposition matrices. However, existing methods perform extensive matrix multiplications in domain specialization tasks, resulting in computational inefficiency and sub-optimal fine-tuning performance. Hence, we propose LoSiA [^1] The source code is released at https://github.com/KlozeWang/LoSiA. (Low-Resources Subnet Integration Adaptation), an innovative method that dynamically localizes and optimizes critical parameters during the training process. Specifically, it identifies a sub-network using gradient sparsity analysis and optimizes it as the trainable target. This design enables effective high-rank adaptation by updating only the sub-network parameters, reducing the additional matrix multiplication. We also present LoSiA-Pro, a faster implementation of LoSiA, which reduces the training latency by about $27\%$ compared to LoRA. Extensive evaluations show that our method achieves minimal performance drop compared to full fine-tuning, while requiring the least training time across domain specialization and common-sense reasoning tasks. Further analysis shows that LoSiA also reduces forgetting during continued training.
@@ 1 Introduction
Figure 1: Overview of LoSiA. It locates and optimizes core sub-network in asynchronous periods.
@@
Large language models, when fine-tuned via supervised learning, can be effectively adapted to downstream tasks such as mathematics Shao et al. (2024), programming Hui et al. (2024), and domain knowledge reasoning Wei et al. (2021). Although full parameter fine-tuning often yields the best performance, updating billions of parameters is computationally expensive and resource-intensive. To address this, parameter-efficient fine-tuning (PEFT) updates only a small subset of parameters to reduce GPU memory usage and communication overhead while maintaining performance comparable to full fine-tuning Houlsby et al. (2019); Ding et al. (2023).
@@
Among PEFT approaches, LoRA Hu et al. (2022) has gained widespread adoption by introducing low-rank matrices to approximate full weight updates, enabling competitive performance with significantly reduced computational and economic costs Taori et al. (2023). Variants in the LoRA family further refine the method by biased fine-tuning modules Zhu et al. (2024); Hayou et al. (2024a) or dimensions Meng et al. (2024a) to accelerate convergence and achieve superior performance. However, constrained by the low-rank assumption, these paradigms often struggle to balance model performance and efficiency, particularly in domain-specific tasks Yang et al. (2024); Ghosh et al. (2024) and continual learning scenarios Shuttleworth et al. (2024a). In such settings, low rank configurations (e.g., $8$ or $16$) can lead to performance degradation and underfitting Biderman et al. (2024). Although increasing the rank may mitigate these issues, it introduces additional memory consumption, extensive floating point operations, and risks of overfitting or convergence difficulties Kalajdzievski (2023); Borse et al. (2024). Recent studies have attempted to approximate high-rank updates by accumulating multiple low-rank components. However, these approaches still suffer from issues such as locally low-rank updates Meng et al. (2024b); Lialin et al. (2023) or increased computational complexity Zhao et al. (2024a). Therefore, while the low-rank assumption offers notable improvements in efficiency, it also introduces inherent limitations.
@@
The Lottery Ticket Hypothesis Frankle and Carbin (2019) suggests that dense neural networks contain trainable sub-networks capable of achieving comparable test accuracy. This prompts us to reconsider the route of PEFT and explore an alternative: Can we identify and fine-tune such sub-networks within the backbone model to achieve high-quality adaptation more efficiently?
@@
To answer this question, we propose LoSiA (Low-Resources Subnet Integration Adaptation), a novel PEFT framework that dynamically localizes and optimizes critical sub-networks periodically, as illustrated in Figure 1. LoSiA asynchronously selects a core sub-network for each layer by calculating sensitivity-based importance scores and performing greedy selecting algorithms. Following localization, it fine-tunes the identified sub-network and applies a rewarming learning rate strategy to promote stable and robust training. The design enables real-time high-rank updates without introducing additional matrix multiplication overhead, which reduces training latency and ensures no steep increase in training time for high-rank updating. Additionally, LoSiA does not introduce extra architectural components and only requires an optimizer replacement for seamless deployment. Extensive experiments demonstrate its superior performance among PEFT baselines on domain-specific, commonsense reasoning tasks, while mitigating forgetting in continue learning. We also propose LoSiA-Pro, a more refined equivalent implementation of LoSiA, which significantly reduces the activation storage and computational complexity in backward propagation. LoSiA-Pro speeds up training $1.38\times$ compared to LoRA and $2.68\times$ compared to DoRA.
@@
(1) Innovatively introduces the structure of the sub-network to the field of parameter-efficient fine-tuning. We develop periodic subnet localization, optimization and integration techniques to dynamically capture and adapt task-essential parameters.
(2) We propose LoSiA, a novel PEFT approach that dynamically localizes and optimizes sub-networks to achieve high performance, efficiency, and usability. To further improve practicality, we also introduce a faster variant, LoSiA-Pro.
(3) We conduct extensive evaluations across multiple models and benchmarks. LoSiA outperforms all advanced PEFT baselines on domain-specific and commonsense reasoning tasks, while also accelerating training by $1.15\times$ compared to LoRA. Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of $1.38\times$.
@@ 2 Related Work
Full parameter fine-tuning (FFT) adapts pre-trained models to downstream tasks by updating all model parameters Wei et al. (2022), but it is often computationally expensive. In contrast, parameter-efficient fine-tuning (PEFT) methods update only a small subset of parameters, significantly reducing training costs while still maintaining strong performance. LoRA Hu et al. (2022) approximates parameter updates as the product of low-rank matrices, achieving promising performance for tasks such as instruction tuning Ghosh et al. (2024). Enhanced variants such as PiSSA Meng et al. (2024a) accelerate convergence by prioritizing major singular vectors, while DoRA Liu et al. (2024) improves performance in low-rank by decomposing updates into directional and magnitude components. Other derivatives such as LoRA+ Hayou et al. (2024b), LoRA-GA Wang et al. (2024a) and LoRA-Dash Si et al. (2025) refine the framework by directional or module biased fine-tuning.
However, recent studies Jiang et al. (2024b); Biderman et al. (2024); Ghosh et al. (2024) reveal that the low-rank structure limits effectiveness in knowledge-intensive domains (e.g., mathematics, coding). Advanced solutions adopt novel strategies: 1) Architectural modifications through MoE-based LoRA combinations Zadouri et al. (2023); Li et al. (2024); Wang et al. (2024b) for multitask scenarios; 2) High-rank fine-tuning via accumulated low-rank projections, such as ReLoRA Lialin et al. (2023), MoRA Jiang et al. (2024a) and GaLore Zhao et al. (2024a) to enhance training effectiveness. However, these approaches incur an increase in architectural complexity or a drop in throughput. Rare methods achieve an optimal balance between performance, training latency, and implementation simplicity.
@@
LLM pruning reduces neural network size by eliminating redundant or less critical parameters. Previous work demonstrates that sparse networks can play crucial roles Frankle and Carbin (2019); Yao et al. (2025). Panigrahi et al. (2023) identifies critical parameters in fine-tuned LMs by optimizing masks of grafted models, although such methods require additional training time and data. Alternatively, gradient- and sensitivity-based metrics enable real-time identification of task-aware parameters Molchanov et al. (2019); Sanh et al. (2020); Zhang et al. (2022). Recent advances adapt the techniques to PEFT: Zhang et al. (2023) prunes LoRA trainable parameters, while Feng et al. (2024) applies the approach to continual learning scenarios.
@@ 3 Method
Consider a model $f_{0}:\mathcal{X}\to\mathcal{Y}$ trained on dataset $\mathcal{D}=\{B_{i}\}_{i=1}^{N}$, where each batch $B_{i}=\{(x_{j},y_{j})\}_{j=1}^{M}$ contains $M$ samples. Let $W$ denote the parameters and $\mathcal{L}$ the loss function. The neural network $S_{0}$ in $f_{0}$ can be represented as a tuple comprising input neurons $X_{S_{0}}$, output neurons $Y_{S_{0}}$ and neural connections $W_{X_{S_{0}},Y_{S_{0}}}$, that is, $S_{0}=(X_{S_{0}},Y_{S_{0}},W_{X_{S_{0}},Y_{S_{0}}})$. The notation $f_{0}\xrightarrow[\mathcal{D}]{P_{0}}f$ denotes training model $f_{0}$ on $\mathcal{D}$ over full parameters $P_{0}=\{W_{X_{S_{0}},Y_{S_{0}}}\}$ and produces model $f$. We investigate the following question: Can we efficiently identify a parameter subset $P\subset P_{0}$ of fixed size, such that training $f\xrightarrow[\mathcal{D}]{P}f^{\prime}$ minimizes the loss difference $\Delta\mathcal{L}=|\mathcal{L}(f^{\prime},\mathcal{D})-\mathcal{L}(f,\mathcal{D})|$?
@@ 3.1 Structure of Gradients
Inspired by pruning techniques, we minimize the mean squared error (MSE) $\mathcal{L}_{\text{MSE}}$ between the outputs of models $f_{k},f^{\prime}_{k}$, where both are trained on $f_{k-1}$ with trainable parameter set $P_{0}$ and $P$, respectively. For SGD optimizers, we derive the bound:
@@
For AdamW optimizers, $\mathcal{L}_{MSE}$ admits a similar bound through $\nabla W$ in most cases (Appendix A.1.1). Thus, gradient magnitudes in $P$ provide an upper bound for the approximation error, while prioritizing larger gradients yields tighter bounds. We therefore seek optimal subsets $P$ to capture the parameters with large gradient magnitudes.
Ideally, selecting the $top\text{-}K$ entries of $\nabla W$ would suffice, but storing and fine-tuning sparse matrices compromises efficiency. Instead, we claim that a suitable pattern for $P$ corresponds to the parameters in subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$, i.e., all connections between the input neuron set $X_{S}$ and the output neuron set $Y_{S}$.
@@
To validate this selection paradigm, Figure 2 visualizes the gradient magnitude distributions in LLaMA-2 7B’s proj_v layer: The 32 attention heads exhibit a significant disparity in attention scores, while gradients strongly correlate with output neurons $Y_{S}$. In particular, a consistent subset of input neurons $X_{S}$ (green markers, x-axis) contributes dominantly to all attention heads. The sparse pattern remains consistent in MLP layers (Appendix A.2.1). We therefore focus on fine-tuning subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$ - termed the core subnet - rather than the entire network.
@@ 3.2 Subnet Localization
To localize core subnets efficiently, an ideal algorithm should satisfy three key requirements: 1) Efficiency: no additional data requirements or significant latency. 2) Lightweight: minimal GPU memory overhead. 3) Dynamic Awareness: enable real-time localizations during training. To meet these objectives, the subnet localization process is divided into two stages:
@@
To quantify parameter importance $I(\cdot)$, existing approaches LeCun et al. (1989); Ma et al. (2023) observe the change in loss assuming $W_{k}=0$ for the $k$-th parameter. Adopting second-order Taylor expansion, element-wise $I(\cdot)$ is estimated as:
@@
Furthermore, estimates by single micro-batch may introduce bias by ignoring training dynamics. Sensitivity smoothing and uncertainty quantification Zhang et al. (2022) are used to handle the problem. For the training step $i$, compute an exponential moving average (EMA) $\overline{I}_{i}$ for $I_{i}$ ,and uncertainty $\overline{U}_{i}$ for variation $\Delta I_{i}=|I_{i}-\overline{I}_{i}|$:
@@
where $\beta_{1},\beta_{2}\in(0,1)$ is the EMA factors. We regard $s(\cdot)$ as a appropriate importance assessment. Notably, to obtain gradients $W$ asynchronously, LoSiA use per-layer weight updates Lv et al. (2024), executing the optimization during backpropagation without storing gradients.
@@
For a subnet $S$ selected from origin network $S_{0}=(\{i\}_{i=1}^{n},\{j\}_{j=1}^{m},W)$, define its importance as:
@@
The objective is to identify optimal subset $S$ and maximize $s(S)$, while satisfying the memory constraint $\max\{\frac{|X_{S}|}{n},\frac{|Y_{S}|}{m}\}\leq p$, where $p\in(0,1)$ represents the rank factor. However, the task is NP-Hard. Leveraging observations in Section 3.1 regarding gradient magnitude sparsity patterns, we develop greedy selection algorithms to identify critical input and output neuron sets $(X_{S},Y_{S})$:
@@
1:function Row2Column($q$)
2:  $\text{sums}\leftarrow\text{Sum}(q,\text{dim}=1)$
3:  $\text{rows}\leftarrow\text{Top-K}(\text{sums},\lfloor np\rfloor).\text{indices}$
4:  $\text{sums}\leftarrow\text{Sum}(q[\text{rows},:],\text{dim}=0)$
5:  $\text{cols}\leftarrow\text{Top-K}(\text{sums},\lfloor mp\rfloor).\text{indices}$
6:  return $(\text{rows},\text{cols})$
@@
Algorithm 1 implements a row-major greedy approach that selects the $top\text{-}K$ rows by importance score summation, and then finds $\{Y_{S}\}$ to maximum $s(S)$ given fixed $\{X_{S}\}$. A column-major greedy algorithm Column2Row is also considered. The better selection result of two strategies is selected.
@@
While prior work Chen et al. (2024) has established the benefits of fine-tuning the output layer in conjunction with PEFT methods, the approach remains computationally prohibitive for large-vocabulary models (e.g., Gemma-2B). However, empirically, backward propagation through the output layer exhibits gradient sparsity, with only a limited subset of tokens receiving significant updates. Building on this insight, LoSiA easily implements an efficient optimization strategy by constructing a tunable subnet $S=(X_{S_{0}},Y_{S},W_{X_{S_{0}},Y_{S}})$ of the output layer, where $|Y_{S}|=p_{o}|Y_{S_{0}}|$, and $p_{o}\in(0,1)$ denoting the dimension reduction factor.

### v4

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA, significantly reduce the number of trainable parameters by introducing low-rank decomposition matrices. However, existing methods perform extensive matrix multiplications in domain specialization tasks, resulting in computational inefficiency and sub-optimal fine-tuning performance. Hence, we propose LoSiA [^1] The source code is released at https://github.com/KlozeWang/LoSiA. (Low-Resources Subnet Integration Adaptation), an innovative method that dynamically localizes and optimizes critical parameters during the training process. Specifically, it identifies a sub-network using gradient sparsity analysis and optimizes it as the trainable target. This design enables effective high-rank adaptation by updating only the sub-network parameters, reducing the additional matrix multiplication. We also present LoSiA-Pro, a faster implementation of LoSiA, which reduces training latency by about $27\%$ compared to LoRA. Extensive evaluations show that our method achieves minimal performance drop compared to full fine-tuning, while requiring the least training time across domain specialization and common-sense reasoning tasks. Further analysis shows that LoSiA also reduces forgetting during continued training.
@@ 1 Introduction
Figure 1: Overview of LoSiA. The method locates and optimizes core sub-network in asynchronous periods.
@@
Large language models, when fine-tuned via supervised learning, can be effectively adapted to downstream tasks such as mathematics Shao et al. (2024), programming Hui et al. (2024), and domain knowledge reasoning Wei et al. (2021). Although full parameter fine-tuning often yields the best performance, updating billions of parameters is computationally expensive and resource intensive. To address this, parameter-efficient fine-tuning (PEFT) updates only small amount of parameters to reduce GPU memory usage and communication overhead while maintaining performance comparable to full fine-tuning Houlsby et al. (2019); Ding et al. (2023).
@@
Among PEFT approaches, LoRA Hu et al. (2022) has gained widespread adoption by introducing low-rank matrices to approximate full weight updates, producing competitive performance with significantly reduced computational and economic costs Taori et al. (2023). Variants in the LoRA family further refine the method by biased fine-tuning modules Zhu et al. (2024); Hayou et al. (2024a) or dimensions Meng et al. (2024a) to accelerate convergence and achieve superior performance. However, constrained by the low-rank assumption, these paradigms often struggle to balance model performance and efficiency, particularly in domain-specific tasks Yang et al. (2024); Ghosh et al. (2024) and continual learning scenarios Shuttleworth et al. (2024a). In such settings, low-rank configurations (e.g., $8$ or $16$) can lead to performance degradation and under-fitting Biderman et al. (2024). Although increasing the rank may mitigate these issues, it introduces additional memory consumption, extensive floating point operations, and risks of overfitting or convergence difficulties Kalajdzievski (2023); Borse et al. (2024). Recent studies have attempted to approximate high-rank updates by accumulating multiple low-rank components. However, these approaches still suffer from issues such as locally low-rank updates Lialin et al. (2023); Meng et al. (2024b) or increased computational complexity Zhao et al. (2024a). Therefore, while the low-rank assumption offers notable improvements in parameter efficiency, it also introduces inherent limitations.
@@
The Lottery Ticket Hypothesis Frankle and Carbin (2019) suggests that dense neural networks contain trainable sub-networks capable of achieving comparable test accuracy. This prompts us to reconsider the PEFT roadmap and explore an alternative: Can we identify and fine-tune such sub-networks within the backbone model to achieve high-quality adaptation more efficiently?
@@
To answer this question, we propose LoSiA (Low-Resources Subnet Integration Adaptation), a novel PEFT framework that dynamically localizes and optimizes critical sub-networks periodically, as illustrated in Figure 1. LoSiA asynchronously selects a core sub-network for each layer by calculating sensitivity-based importance scores and performing greedy selecting algorithms. Following localization, it fine-tunes the identified sub-network and applies a rewarming learning rate strategy to promote stable and robust training. The design enables real-time high-rank updates without introducing additional matrix multiplication overhead, while significantly reducing training latency. Additionally, LoSiA does not introduce extra architectural components and only requires optimizer replacements for seamless deployment. Extensive experiments demonstrate its superior performance among PEFT baselines on domain-specific, commonsense reasoning tasks, while mitigating forgetting in continue learning. We also propose LoSiA-Pro, a more refined equivalent implementation of LoSiA, which significantly reduces the activation storage and computational complexity in backward propagation. LoSiA-Pro speeds up training $1.38\times$ compared to LoRA and $2.68\times$ compared to DoRA.
@@
(1) Innovatively, we incorporate sub-network structure into the field of parameter-efficient fine-tuning. We devise a periodic workflow with techniques that localize, optimize, and integrate sub-networks, thus flexibly capturing and adapting task-essential parameters.
(2) We propose LoSiA, a novel high performance PEFT approach that dynamically localizes and optimizes sub-networks. By eliminating redundant computation, we further propose LoSiA-Pro, a loss-less variant that markedly reduces training latency and GPU memory footprint.
(3) We conduct extensive evaluations across multiple models and benchmarks. LoSiA outperforms all advanced PEFT baselines on domain-specific and common-sense reasoning tasks, while also accelerating training $1.15\times$ compared to LoRA. Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of $1.38\times$.
@@ 2 Related Work
Full param-eter fine-tuning (FFT) adapts pre-trained models to downstream tasks by updating all model parameters Wei et al. (2022), yet incurs prohibitive computational overhead. In contrast, parameter-efficient fine-tuning (PEFT) methods update only a small subset of parameters, curbing training costs while sustaining competitive accuracy. LoRA Hu et al. (2022) approximates parameter updates as the product of low-rank matrices, achieving promising performance in tasks such as instruction tuning Ghosh et al. (2024). Enhanced variants such as PiSSA Meng et al. (2024a) accelerate convergence by prioritizing dominant singular vectors, while DoRA Liu et al. (2024) decomposes updates into directional and magnitude components for more effective fine-tuning. Other derivatives such as LoRA+ Hayou et al. (2024b), LoRA-GA Wang et al. (2024a), and LoRA-Dash Si et al. (2025) refine the framework by directional or module biased optimization.
However, recent studies Jiang et al. (2024b); Biderman et al. (2024); Ghosh et al. (2024) reveal that the low-rank bottleneck restricts effectiveness in knowledge-intensive domains (e.g., mathematics, coding). Advanced solutions adopt strategies such as: 1) Architectural modifications through MoE-based LoRA combinations Zadouri et al. (2023); Li et al. (2024); Wang et al. (2024b) for multitask learning scenarios; 2) High-rank fine-tuning via accumulated low-rank projections, such as ReLoRA Lialin et al. (2023), MoRA Jiang et al. (2024a) and GaLore Zhao et al. (2024a) to enhance training effectiveness. However, these ameliorated approaches either inflate architectural complexity or compromise throughput. Rare methods simultaneously optimize performance, training latency, and implementation simplicity.
@@
LLM pruning compresses networks by excising redundant or less critical parameters. Previous work demonstrates that sparse networks can play crucial roles Frankle and Carbin (2019); Yao et al. (2025). Panigrahi et al. (2023) identifies critical parameters in fine-tuned LMs by optimizing masks of grafted models, but such methods require additional training time and data. Alternatively, gradient- and sensitivity-based metrics enable real-time identification of task-aware parameters Molchanov et al. (2019); Sanh et al. (2020); Zhang et al. (2022). Recent advances extend these ideas to PEFT: Zhang et al. (2023) prunes LoRA trainable parameters, while KIF (Feng et al., 2024a, b) integrates skill localization into continual-learning regimes.
@@ 3 Method
Consider a model $f_{0}:\mathcal{X}\to\mathcal{Y}$ trained on dataset $\mathcal{D}=\{B_{i}\}_{i=1}^{N}$, where each batch $B_{i}=\{(x_{j},y_{j})\}_{j=1}^{M}$ contains $M$ samples. Let $W$ denote the parameters and $\mathcal{L}$ the loss function. The neural sub-network $S$ in $f_{0}$ is represented as the tuple $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$, comprising its input neurons $X_{S}$, output neurons $Y_{S}$ and neural connections $W_{X,Y}$. Training model $f_{0}$ on $\mathcal{D}$ with full parameters $P_{0}=W$ is compactly written as $f_{0}\xrightarrow[\mathcal{D}]{P_{0}}f$. We investigate the following question: Given a cardinality budget, can we efficiently identify a parameter subset $P\subset P_{0}$, such that $f_{0}\xrightarrow[\mathcal{D}]{P}f^{\prime}$ minimizes the loss difference $\Delta\mathcal{L}=|\mathcal{L}(f^{\prime},\mathcal{D})-\mathcal{L}(f,\mathcal{D})|$?
@@ 3.1 Structure of Gradients
Inspired by pruning techniques, we expect to minimize the mean squared error (MSE) $\mathcal{L}_{\text{MSE}}$ between subsequent models $f_{k}$ and $f^{\prime}_{k}$, where they both are trained from the model $f_{k-1}$ with trainable parameters $P_{0}$ and $P$, respectively. For SGD optimizers, we derive the bound:
@@
For AdamW, an analogous bound $\mathcal{L}_{MSE}$ holds in terms of $\nabla W$ in most cases (Appendix A.1.1). Thus, gradient magnitudes lay in $P$ provide an upper bound for the approximation error, while retaining parameters with the largest gradient magnitudes to adjust tightens the bound.
Ideally, selecting the $\text{Top-}K$ entries of $\nabla W$ is theoretically optimal, but storing and fine-tuning sparse matrices compromises the efficiency. Instead, we claim that a suitable selection pattern for $P$ corresponds to a structured subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$, i.e., all connections between input neuron set $X_{S}$ and output neuron set $Y_{S}$.
@@
To validate this selection paradigm, Figure 2 visualizes the gradient magnitude distributions in LLaMA-2 7B’s proj_v layer. Across the 32 attention heads, the gradient norms exhibit pronounced skewness and are highly correlated with the corresponding output neurons $Y_{S}$. On the other hand, a consistent set of input neurons $X_{S}$ (green markers, x-axis) contributes dominantly to all attention heads. The sparse pattern also holds in MLP layers (Appendix A.2.1). Consequently, we restrict the fine-tuning space to subnet structures $S$ - termed the core subnet - rather than the entire network.
@@ 3.2 Subnet Localization
To efficiently localize core subnets, an ideal algorithm should satisfy three key requirements: 1) Efficiency: no extra data or heavy computation. 2) Lightweight: negligible GPU memory overhead. 3) Dynamic Awareness: enable on-the-fly localization throughout the training process. Although existing LLM-pruning methods have achieved impressive compression ratios, they still fall short of simultaneously satisfying the aforementioned desiderata. Consequently, we devise a dedicated subnet-localization algorithm tailored for efficient fine-tuning that is divided into two stages:
@@
To quantify parameter importance $I(\cdot)$, existing approaches LeCun et al. (1989); Ma et al. (2023) observe the change in loss by assuming $W_{k}=0$ for the $k$-th parameter. Adopting the second-order Taylor expansion, element-wise importance is estimated as:
@@
Furthermore, estimation with single micro-batch may inject bias by overlooking training dynamics. Sensitivity smoothing and uncertainty quantification Zhang et al. (2022) are used to handle the problem. At training step $i$, it maintains an exponential moving average (EMA) $\overline{I}_{i}$ for $I_{i}$, and uncertainty $\overline{U}_{i}$ for variation $\Delta I_{i}=I_{i}-\overline{I}_{i}$:
@@
where $\beta_{1},\beta_{2}\in(0,1)$ are the EMA factors. We treat $s(\cdot)$ as an appropriate importance assessment. To obtain the weight-gradient signal without keeping all full tensors in memory, LoSiA uses per-layer updates Lv et al. (2024), executing optimizations during backpropagation without storing gradients.
@@
Given a subnet $S$ of the origin network $S_{0}=(\{i\}_{i=1}^{n},\{j\}_{j=1}^{m},W)$, define its importance as:
@@
Our objective is to identify the optimal subnet $S$ that maximizes $s(S)$, while respecting the memory cap $\max\{\frac{|X_{S}|}{n},\frac{|Y_{S}|}{m}\}\leq p$, where $p\in(0,1]$ represents the rank factor. However, the task is NP-Hard. Exploiting the gradient-magnitude sparsity patterns observed in Section 3.1, we develop greedy selection algorithms to select the critical input and output neuron set $X_{S}$ and $Y_{S}$.
@@
1:Importance matrix $s\in\mathbb{R}^{n\times m}$ with $s_{ij}=s(W_{ij})$; rank factor $p\in(0,1]$.
[…]
3:function Row2Column($s\in\mathbb{R}^{n\times m},\ p$)
4:  $\rho\leftarrow\operatorname{Top-K\ Indices}\!\ \bigl(\sum_{j=1}^{m}s_{:,j},\;\lfloor np\rfloor\bigr)$
5:  $\gamma\leftarrow\operatorname{Top-K\ Indices}\!\ \bigl(\sum_{i\in\rho}s_{i,:},\;\lfloor mp\rfloor\bigr)$
6:  return $(\rho,\gamma)$
@@
Algorithm 1 embodies a row-major greedy policy. First, it locks the $\lfloor np\rfloor$ input neurons with the highest row-wise aggregate importance, then greedily retains the $\lfloor mp\rfloor$ output neurons that maximize the residual mass in those fixed rows. A symmetric column-major variant reverses the order of fixation. The final subnet adopts whichever of the two masks yields the higher score $s(S)$.
@@
Although prior work Chen et al. (2024) has established the benefits of fine-tuning the output layer in conjunction with PEFT methods, the approach remains computationally prohibitive for large-vocabulary models (e.g. Gemma-2B). However, empirically, backward propagation through the output layer exhibits gradient sparsity, with only a limited subset of tokens receiving significant updates. Building on this insight, LoSiA easily implements an efficient optimization strategy by constructing a tunable subnet $S=(X_{S_{0}},Y_{S},W_{X_{S_{0}},Y_{S}})$ in the output layer, where $|Y_{S}|=p_{o}|Y_{S_{0}}|$, and $p_{o}\in(0,1]$ denote the output dimension reduction factor.